# Explainable Machine Learning for ICU Readmission Prediction

### v1

@@ Abstract
The intensive care unit (ICU) comprises a complex hospital environment, where decisions made by clinicians have a high level of risk for the patients’ lives. A comprehensive care pathway must then be followed to reduce patient complications. Uncertain, competing and unplanned aspects within this environment increase the overall difficulty in uniformly implementing the care pathway. Readmission contributes to this pathway’s difficulty, occurring when patients are admitted again to the ICU in a short timeframe, resulting in high mortality rates and high resource utilisation. Several works have tried to predict readmission through patients’ medical information. Although they have some level of success while predicting readmission, those works do not properly assess, characterise and understand readmission prediction. This work proposes a standardised and explainable machine learning pipeline to model patient readmission on a multicentric database (i.e., the eICU cohort with 166,355 patients, 200,859 admissions and 6,021 readmissions) while validating it on monocentric (i.e., the MIMIC IV cohort with 382,278 patients, 523,740 admissions and 5,984 readmissions) and multicentric settings. Our machine learning pipeline achieved predictive performance in terms of the area of the receiver operating characteristic curve (AUC) up to 0.7 with a Random Forest classification model, yielding an overall good calibration and consistency on validation sets. From explanations provided by the constructed models, we could also derive a set of insightful conclusions, primarily on variables related to vital signs and blood tests (e.g., albumin, blood urea nitrogen and hemoglobin levels), demographics (e.g., age, and admission height and weight), and ICU-associated variables (e.g., unit type). These insights provide an invaluable source of information during clinicians’ decision-making while discharging ICU patients.
@@
[label14]organization= The Department of Intensive Care, Ballarat Base Hospita, city=Ballarat, state=Victoria, country=Australia
@@ 1 Introduction
Intensive care is a service for patients with likely recoverable and treatable conditions, providing them with more precise observation and invasive treatments Smith and Nielsen (1999). Appropriate patient management within an intensive care unit (ICU) is crucial. In a given ICU, the likelihood of long-term length of stay (LOS), organ failures, and mortality tends to increase if adequate management is not taken into consideration Moitra et al. (2016); Sakr et al. (2012); Ay et al. (2020). Clinicians who are caring for these patients also have multiple competing activities to consider Sujan et al. (2015); Turnbull et al. (2018), including the potential deterioration of the patient’s condition after discharge, emergency admissions, elective admissions for high-risk surgery, staffing and resources, and evidence for appropriate allocation of resources. These aspects highlight the challenges across an ICU setting.
@@
Readmission is one of the factors that extends the challenges within an ICU setting. It occurs when patients are admitted again to the ICU in a short timeframe (between 48 hours and 30 days), resulting in high mortality rates, increased LOS and, consequently, high resource utilisation Mcneill et al. (2020); Alban et al. (2006); Kramer et al. (2013). In summary, when readmission occurs, it disrupts the care pathway for the patients and poses additional challenges to the clinical team and hospital caring for them.
@@
Reducing ICU readmission rates might not only improve the care pathway, leading to better patient outcomes but also affect the hospital’s bottom line James (2013); Khera and Krumholz (2018). Predicting patients at high risk of readmission would not only allow early intervention to reduce the risk but also reduce mortality rates, reduce resource utilisation (based on LOS) and, potentially, hospital costs.
@@
The emergence of abundant hospital electronic health record (EHR) data Yadav et al. (2018); Kim et al. (2019); Shickel et al. (2017) allowed the use of machine learning (ML) models targeting ICU readmission Badawi and Breslow (2012); Rojas et al. (2018); Mohanty et al. (2022); Thoral et al. (2021); McWilliams et al. (2019); Hammer et al. (2021); Lin et al. (2019); Xue et al. (2019); Barbieri et al. (2020); Hegselmann et al. (2022); Ashfaq et al. (2019). These readmission models have been summarised in Table S1. Apart from their undeniable importance, these models lack proper interpretation, generalisation and validation in different ICU settings (e.g., monocentric versus multicentric environments).
@@
Our main contention is that predicting ICU readmission in a generalisable way with hospital-based datasets is still a challenge to overcome, even after taking into account these univariate predictive models and best practice ML techniques Lin et al. (2019); Hosein et al. (2013); Curto et al. (2016). The main reason for this difficulty is that medical data is usually noisy, uncertain, and characterised by a high degree of missingness, as it relies on different human inputs. Further exacerbating this complexity is data imbalance, with only a small proportion of patients being readmitted. ICU patients who happen to be readmitted are also very heterogeneous Hosein et al. (2013), meaning that finding novel patterns across an ICU readmission data cohort tends to be challenging. As a result, an appropriate translation from ML models to the clinical environment is still limited.
In addition, readmission’s association with healthcare quality has been debated in recent years Brown et al. (2014); Rewa et al. (2018). Many readmissions also emerge from factors inherent to patients, not only the quality of care, and may be reduced but not entirely eliminated. In these cases, better healthcare quality might not prevent readmission Pakbin et al. (2018). Still, predicting readmission accurately is valuable for post-discharge monitoring and planning. Machine learning models can connect clinical data to readmission, aiding pre-discharge decisions and informing readmission reduction programs Rojas et al. (2018); Thoral et al. (2021); Huang et al. (2021); Hu et al. (2022); McIlvennan et al. (2015).
This work proposes a standardised method for a comprehensive assessment of readmission modelling and prediction on ICU data Pollard et al. (2018, 2019) with explainable machine learning Molnar (2023). Our method resulted in a generalisable predictive model for both multicentric and monocentric data settings Pollard et al. (2018, 2019); Johnson et al. (2021). This is the first study to provide such a large and robust validation on readmission models. In addition, explanations from the constructed model also provided a better understanding of the main relevant patient variables associated with 30-day ICU readmission, which can potentially help in guiding clinicians while treating and discharging patients from the ICU.
1.1 State of Significance
Problem Intensive Care Unit (ICU) Readmission Prediction with Explainable Machine Learning.
What is Already Known Although there has been a range of related works aiming to predict ICU readmission with a variable level of success, those studies do not properly assess, characterise and understand readmission prediction, limiting their overall applicability to patient data.
What This Paper Adds This study proposes a standardised and explainable machine learning (ML) pipeline to model, predict and understand 30-day readmission. With a robust validation procedure, the developed ML model achieved generalisable performance on completely new patient data while deriving insightful (literature-supported) conclusions to contrast readmission and non-readmission within an ICU.
@@ 2 Material and Methods
Figure 1 presents the main methodological workflow followed in this work. Collecting data from monocentric (i.e., MIMIC IV) and multicentric databases (i.e., eICU) presents itself as the first step. Next, from both databases, the creation of variables describing the patient’s information takes place considering their admission timeframe. These variables include patient demographics, pre-admission information, vital signs and blood test results, and take into account the first and the last 24 hours, as well as their average values. Data preprocessing includes several steps, such as imputation, imbalance learning, and feature standardisation. The inclusion of variables outlining the hospitals and ICUs also plays an important role in this characterisation. Machine learning algorithms take these features as input to build a readmission model based on a timeframe of 30 days. The employed machine learning algorithms use multicentric data from eICU to build and internally validate the predictive model for readmission while it externally tests and validates the final model on monocentric data from MIMIC-IV. The readmission model predicts whether a given patient will be readmitted or not and supports explanations for its predictions, driving insights to clinicians while discharging a given patient in the ICU.
@@
Figure 1: The methodological workflow followed by this work. Data comes from monocentric and multicentric databases (i.e., MIMIC IV and eICU, respectively), where the characterisation of ICU patients, hospitals, and ICU takes place. Preprocessing filters out, standardises and imputes data for the development of a 30-day readmission machine learning model, which is built and validated on eICU, and tested on MIMIC-IV. This model drives explanations from variables and predictions, potentially assisting and guiding clinicians while treating and discharging new patients from the ICU.
2.1 Data Sources
For assessing, analysing and understanding ICU readmission, we utilise two datasets: (i) the eICU Collaborative Research Database (eICU) Pollard et al. (2018, 2019); Goldberger et al. (2000), and (ii) the Medical Information Mart for Intensive Care (MIMIC IV) Johnson et al. (2021); Goldberger et al. (2000), which are described in next subsections. Their access relied on PhysioNet Goldberger et al. (2000).
2.1.1 eICU
The eICU contains data collected from 335 ICUs at 208 US hospitals in 2014 and 2015. This cohort has information about 166,355 unique patients and 200,859 admissions, accounting also for readmissions for those listed patients.
The eICU cohort is a de-identified multicentric dataset with information about patient demographics, admission and discharge details, APACHE IV scores and variables used for its calculation, vital signs, laboratory test results, medications, and care plan information.
2.1.2 MIMIC IV
The MIMIC cohort (version IV) emerged from two in-hospital systems of the Beth Israel Deaconess Medical Center (BIDMC): Electronic Health Record (EHR) and ICU-specific Clinical Information System. MIMIC IV has data about patients’ admissions through the emergency department or one of the intensive care units at BIDMC between 2008 and 2019. It includes data about 382,278 patients and 523,740 admission records.
Like eICU, this database is also de-identified and comprises demographics, hospital billing information, data related to patients’ admissions, transfers, and discharges, information about medication administration, and laboratory measurements. MIMIC IV also contains data generated in the ICU, such as intravenous and fluid inputs, procedures and services, charted information, and patients’ outputs.
2.2 Readmission Outcome
Patients were considered to have met the primary readmission outcome if they had one or more admission episodes within 30 days of discharge from their first ICU admission (see Supplementary Material and Methods for more details). If a patient is readmitted twice or more times, only the first readmission within 30 days counts for their data. It is worth noting that this range of days to define readmission follows different methodologies Mohanty et al. (2022); Lin et al. (2019); Xue et al. (2019); Barbieri et al. (2020).
Within the 30 days, the eICU cohort had a total of 149,009 patients, where 6,021 and 142,988 were admitted and non-readmitted, respectively. On the other hand, MIMIC IV had a total of 68,505 patients, where 5,984 underwent readmission and 62,521 did not.
2.3 Predictor Variable Selection
The same potential predictor variables were extracted from both datasets, which resulted in 168 common variables. Patient-level data included demographics (age, sex, weight, and height), admission diagnosis, highest and lowest physiological and biochemical values from the first and last 24 hours of ICU admission, treatments provided in ICU (ventilation, dialysis, and vasopressor) and overall illness severity scores (i.e., APACHE IV scores Lee et al. (2015)). Institutional-level variables included hospital-related (e.g., hospital size and location) and ICU-related variables (e.g., unit type and admission source). Supplementary Materials (Table S2) detail and distinguish these common variables in both databases across the following six categories: (i) demographics, (ii) vital sign- and blood test-related, (iii) pathology- and treatment-based, (iv) APACHE severity-related, (v) hospital-related, and (vi) ICU-related.
In addition, Figure S1 and Supplementary Material and Methods cover more details regarding the data collection criteria and the prediction window. Furthermore, categorical variables were encoded in a one-hot encoding scheme (a.k.a, one-of-K or dummy). Given the transformation in the categorical feature space, the number of features increased from 168 to 186 variables.
It is also worth noting that eICU contains an extra level of information on variables when compared to MIMIC IV. The variables from the aforementioned categories, which are detailed in Table S2, are only those that match across both cohorts or do not have linkage to the target readmission variable (e.g., length of stay or mortality). Nevertheless, as we used eICU to train and validate the machine learning models, we have included important eICU variables (missing in MIMIC IV) to analyse readmission. This step has been done only for the sake of readmission explainability. This list of complementary variables is detailed in Table S3.
2.4 Data Preprocessing
We performed a standardised approach to preprocess both datasets (i.e., eICU and MIMIC IV). This approach included basic steps to filter patients’ information and yield valid and reliable information for readmission prediction.
2.4.1 General Filtering
First, this step removed patients that are younger than 18 years old from the cohort. In addition, the proposed filter did not consider patients who died during their first ICU. Furthermore, it separated patients who spent less than four hours in the ICU from the datasets. This preprocessing step also excluded patients with abnormal values in the generated variables, such as having an admission weight greater than 260 kg or having an admission height greater than 240 cm. Supplementary Material and Methods and Figure S2 detail these patients’ inclusion and exclusion criteria on both eICU and MIMIC IV cohorts.
2.4.2 Handling Missing Data
The employed imputation method took into consideration a threshold of 20% of missingness, meaning that our machine learning pipeline will only use variables with less than 20% of missing values. The imputation method handled missing values by replacing them with population means (in the case of continuous variables) or a constant (in the case of categorical variables).
2.4.3 Dealing with Imbalance
Given the low proportion of readmitted patients in the ICU, the proposed pipeline considered several methods from the Imbalanced-learn toolbox Lemaître et al. (2017) to tackle the hardness of learning on imbalanced data Fernández et al. (2018). Nevertheless, random undersampling over the majority class (i.e., the non-readmitted class) yielded similar results to other methods, being then selected by its simplicity and performance.
Random undersampling took into accounting for the different data distributions, from 1:1 (1 readmitted patient to 1 non-readmitted patient) to 1:9 (one readmitted patient to 9 non-readmitted patients) to the real distribution (approximately one readmitted patient to 25 non-readmitted patients in eICU, and one readmitted patient to 10 non-readmitted patients in MIMIC IV).
2.4.4 Standardisation
In keeping with best practices to improve the predictive performance of variables Raschka et al. (2022), standardisation was undertaken by transforming each value of a given variable. On each variable value, standardisation subtracts it from the variable’s mean and then divides it by the variable’s standard deviation. Although standardisation is not a matter for some machine learning models, it is for several others. We therefore applied the same pipeline for all of them, making sure the use of proper data Raschka et al. (2022).
2.5 Machine Learning Modelling, Evaluation and Feature Selection
Machine learning (ML) modelling followed traditional development and evaluation guidelines Raschka (2018). The proposed modelling included the use of seven ensemble-based ML algorithms for patient readmission prediction: (i) explainable boosting machine (EBM) Nori et al. (2019), (ii) extremely randomised trees (ERT) Geurts et al. (2006), (iii) random forest (RF) Breiman (2001), (iv) extreme gradient boosting (XGBOOST) Chen and Guestrin (2016), (v) adaptive boosting (ADABOOST) Freund and Schapire (1995), (vi) gradient boosting (GB) Friedman (2001), and (vii) light gradient boosting machine (LightGBM) Machado et al. (2019). Table S4 presents the hyper-parameter optimisations for these ML algorithms.
It is important to emphasise that these algorithms have been chosen not only because of their robustness based on ensemble properties (i.e., by combining the results of several ML models) but also for their interpretation easiness. They all have (at some level) pre-built functions to describe feature importance for a given assessed dataset and can be straightforwardly used with other methods/tools for explainability.
With these algorithms as base classifiers, a bottom-up greedy feature selection ensued to ensure low ML complexity while maintaining generalisable predictive performance. Greedy feature selection starts its iterative process by having a bucket with zero features, adding one feature at each round. This method adds a feature in its bucket if it achieves the best predictive performance on a 10-fold cross-validation procedure, which relies on a model built by a classification algorithm. The evaluation of each feature relies on an average between Matthew’s correlation coefficient (MCC) Chicco and Jurman (2020) and the area under the receiver operating characteristic (AUC) curve Raschka et al. (2022); Raschka (2018); Sokolova and Lapalme (2009), which therefore considers resilience to class imbalance. In addition, other evaluation measures complemented the decision process in our ML pipeline, including the area under the precision-recall curve (APR), balanced accuracy, precision, recall (sensitivity), specificity, and F1 score Raschka et al. (2022); Raschka (2018); Sokolova and Lapalme (2009).
2.6 Statistical Analyses
Data are presented as number/proportion (percentage), mean (± standard deviation) or median (inter-quartile range) depending on the type and distribution of data. Chi-squared, ANOVA, Kruskal-Wallis and Wilcoxon rank-sum tests were used to compare groups depending on the type and distribution of data and the number of groups examined Demšar (2006). Statistical significance considers p-values lower than 0.001. The sample size of eICU and MIMIC IV datasets determined this decision, which consequently guaranteed the robustness of this study.
It is also worth noting that statistical analyses performed readmission comparisons not only within the same population (e.g., eICU) but also between populations (i.e., eICU versus MIMIC IV), strengthening this methodology. Furthermore, APACHE scores are provided only for eICU databases, while not directly available for MIMIC IV, enriching the explanation of characteristics of the former to the latter.
2.7 Model’s Diagnostic Interpretation, Analysis, and Explanations
The proposed ML pipeline for readmission also assesses its final model with a posteriori interpretation and explanation methods, which are described next.
2.7.1 Likelihood Ratios
Likelihood ratio (LR) McGee (2002) provides a measure for expressing the diagnostic accuracy of a clinical finding of the readmission model from this work. LR represents the probability of patients with a condition or undergoing a medical event (e.g., being readmitted) divided by the probability of patients without a condition or undergoing a medical event (e.g., not being readmitted). Positive LRs in readmission prediction estimate the model’s capabilities of identifying readmitted against non-readmitted patients. When LRs are greater than one, it means they go along with the readmission diagnosis. In other words, the greater the LR is, the more certain the predictive model is in finding readmitted patients against non-readmitted patients.
2.7.2 Calibration
The calibration technique visually evaluates the predictive readmission model by inspecting the calibration curves and also numerically estimating the calibration slope and intercept, as well as the integrated calibration index (ICI) and associated metrics (E50, E90, Emax ) Austin and Steyerberg (2019). The ideal calibration slope is equal to one. The ideal calibration intercept, as well as the ideal value for the ICI and its related metrics, is zero.
2.7.3 Shapley Additive Explanations
Our pipeline also utilised Shapley Additive Explanations (SHAP) Lundberg and Lee (2017), which is a game theory-based method to derive explanations for individual predictions Molnar (2023), in which combined results yield a global interpretation of the model’s behaviour. In readmission prediction, SHAP assists with model analysis, leading to a better understanding of readmission through a range of explanations. As a result, it can be used to recommend a set of personalised ICU guidelines while treating/discharging patients in the ICU.
@@ 3 Results
Apart from that, we could observe they have similar baseline characteristics (Table 2). However, MIMIC IV readmitted patients received more ventilation and vasopressors but fewer renal replacement therapies than the readmitted eICU patients. In addition, MIMIC IV readmitted patients stayed longer in the ICU and hospital stay. Mortality was also slightly higher in the MIMIC IV readmission patients, although no statistical difference has been found between the two databases. Among survivors, the readmission group in the eICU was discharged more to home, while readmitted patients in MIMIC IV had a higher rate of further treatments in a rehabilitation care facility (p-value <<< 1%).

### v4

@@ Abstract
Background
The intensive care unit (ICU) comprises a complex hospital environment, where decisions made by clinicians have a high level of risk for the patients’ lives. Uncertain, competing and unplanned aspects within the ICU environment increase the difficulty in uniformly implementing the care pathway. Readmission contributes to this pathway’s difficulty, resulting in high mortality rates and resource utilisation.
Objectives
Several works have tried to predict readmission through patients’ medical information. Although they have some level of success while predicting readmission, those works do not properly assess, characterise and understand readmission prediction. This work proposes a standardised and explainable machine learning pipeline to model, predict and explain patient ICU readmission.
Method
This work focuses on using machine learning to model readmission using a multicentric eICU database while validating it on both MIMIC-IV monocentric and eICU multicentric settings. SHAP (Shapley Additive Explanations), calibration and likelihood ratio analysis are employed to comprehend the behaviour of the predictive model.
Results
Our machine learning pipeline achieved predictive performance in the area of the receiver operating characteristic curve (AUC) up to 0.7 with a Random Forest classification model, yielding an overall consistency and generalisation on validation sets. After an interpretation analysis of our model through likelihood ratios and calibration, we observed that the proposed methodology could generate predictive models with proper and translatable diagnosis capabilities. In addition, from explanations provided by the constructed model, we could also derive a set of insightful conclusions, primarily on variables related to vital signs and blood tests (e.g., albumin, blood urea nitrogen and haemoglobin levels), demographics (e.g., age, and admission height and weight), and ICU-associated variables (e.g., unit type).
Conclusions
The model’s predictions, behaviour and explainability method provided insights that yielded valuable information, which clinicians might (be able to) use to make decisions while discharging ICU patients.
@@ Conclusions
[label14]organization=The Department of Intensive Care, Ballarat Base Hospital, city=Ballarat, state=Victoria, country=Australia
@@ 1 Introduction
Intensive care is a service for patients with likely recoverable and treatable conditions, providing them with more precise observation and invasive treatments [1]. Appropriate patient management within an intensive care unit (ICU) is crucial. In a given ICU, the likelihood of long-term length of stay (LOS), organ failures, and mortality tends to increase if adequate management is not taken into consideration [2, 3, 4, 5]. Clinicians who are caring for these patients also have multiple competing activities to consider [6, 7], including the potential deterioration of the patient’s condition after discharge, emergency admissions, elective admissions for high-risk surgery, staffing and resources, and evidence for appropriate allocation of resources. These aspects highlight the challenges across an ICU setting.
@@
Readmission is one of the factors that extends the challenges within an ICU setting. It occurs when patients are admitted again to the ICU in a short timeframe (between 48 hours and 30 days), resulting in high mortality rates, increased LOS and, consequently, high resource utilisation [5, 8, 9, 10]. In summary, when readmission occurs, it disrupts the care pathway for the patients and poses additional challenges to the clinical team and hospital caring for them.
@@
Reducing ICU readmission rates might not only improve the care pathway, leading to better patient outcomes, but also affect the hospital’s bottom line [11, 12]. Predicting patients at high risk of readmission would not only allow early intervention to reduce the risk but also reduce mortality rates, reduce resource utilisation (based on LOS) and, potentially, hospital costs.
@@
The emergence of abundant hospital electronic health record (EHR) data [13, 14, 15] allowed the use of machine learning (ML) models targeting ICU readmission [16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27]. These models for general readmission prediction have been summarised in Table S1. Apart from their undeniable importance, these models lack proper interpretation, generalisation and validation in different ICU settings – e.g., monocentric (one hospital centre) versus multicentric (multiple ICU hospital centres).
@@
Our main contention is that predicting ICU readmission in a well-generali-sable way with hospital-based datasets is still a challenge to overcome, even after taking into account these univariate predictive models and best practice ML techniques [22, 28, 29]. The main reason for this difficulty is that medical data is usually noisy, uncertain, and characterised by a high degree of missingness, as it relies on different human inputs. Further exacerbating this complexity is data imbalance, with only a small proportion of patients being readmitted. ICU patients who happen to be readmitted are also very heterogeneous [28], meaning that finding novel patterns across an ICU readmission data cohort tends to be challenging. As a result, an appropriate translation from ML models to the clinical environment is still limited. To improve the clinical translation, databases (e.g., MIMIC-IV and eICU) should be homogenised and have their noise and missingness diminished.
Accordingly, this work proposes a standardised methodological pipeline for a comprehensive assessment of ICU readmission modelling, prediction and understanding [30, 31] using machine learning models [32, 33] and statistical analyses. The produced predictive ICU readmission model – which was trained and internally validated in a dataset (i.e., eICU) containing several hospitals (multicentric) and externally validated in a monocentric dataset (i.e., MIMIC IV) [30, 31, 34] – shows clear signs of predictive generalisation. It is the first time in the literature that such an approach is employed to confirm the performance of an ICU readmission model. To the best of our knowledge, no other previous method has been validated in opposing ICU settings (i.e., monocentric and multicentric).
@@ 2 Material and Methods
Figure 1 presents the followed methodological workflow. The first step is to collect data from two databases: MIMIC IV (Medical Information Mart for Intensive Care, Version IV) [34, 35] and eICU (eICU Collaborative Research Database) [30, 31, 35]. From these databases, variables describing the patient’s information are created considering their admission timeframe (see Figure S1). These variables include patient demographics, pre-admission information, vital signs and blood test results, and also variables outlining the hospitals and ICUs. These variables are fully described in Tables S2 and S3.
@@
Data preprocessing considers several steps, e.g., imputation, imbalance learning, and feature standardisation. Exclusion criteria are also taken into consideration to avoid abnormal patient information in our dataset (see Figure S2). Machine learning algorithms take these preprocessed features as input to build a readmission model based on a timeframe of 30 days. Hyperparameter optimisation also occurs (Table S4), guaranteeing the selection of the most appropriate predictive model for the data at hand. The final employed machine learning algorithms use multicentric data from eICU to build and internally validate the predictive model for readmission while externally testing and validating the final model on monocentric data from MIMIC-IV. The readmission model predicts whether a given patient will be readmitted or not and supports explanations for its predictions, driving insights to clinicians while discharging a given patient in the ICU.
@@
Figure 1: The methodological workflow followed by this work. Data comes from monocentric and multicentric databases (i.e., MIMIC IV and eICU, respectively), where the characterisation of ICU patients, hospitals, and ICU(s) takes place. Preprocessing filters out, standardises and imputes data for the development of a 30-day readmission machine learning model, which is built and validated on eICU and tested on MIMIC-IV. This model is interpreted and also used to drive explanations from variables and predictions, potentially assisting and guiding clinicians while treating and discharging new patients from the ICU.
The Supplementary Material (more specifically, Supplementary Methods) provides complete details about data sources (eICU, MIMIC IV, and readmission outcome), predictor variables, data preprocessing, machine learning modelling and validation, and the model’s explanation, diagnosis expression, and calibration.
@@ 3 Results
Apart from that, we could observe they have similar baseline characteristics (Table 2). However, MIMIC IV readmitted patients received more ventilation and vasopressors but fewer renal replacement therapies than the readmitted eICU patients. In addition, MIMIC IV readmitted patients stayed longer in the ICU and hospital stay. Mortality was also slightly higher in the MIMIC IV readmission patients, although no statistical difference has been found between the two databases. Among survivors, the readmission group in eICU was discharged more to home, while readmitted patients in MIMIC IV had a higher rate of further treatments in a rehabilitation care facility (p-value <<< 1%).